PS

LIU WS installed. 5 BPMs installed and FT16 MSF in TT2

BGI in SS82, operational WS and BTV MU41 next week.

TT2 limited access.

 

Jeroen wanted to remove 43 pre-amplifiers. He will not have time to do this all.

 

BLMs

6 detectors installed in PSB. Want to install more, but may have to remove them as the documentation is not ready.

 

SPS

Tune pickup mainineane in 2 weeks.

BSI in TT20 - not sure if this will happen. Not much time left.

DBLM is included in another IMPACT so can remove the line in red.

 

LHC

Concern that there is a vacuum problem in the BSRT sector - Gerhard will check this.

BCTDC is under modification - to be checked if this is complete.

Opening of vacuum sector in p7 for crystal collimation.

Christos says that CMS are asking for valve opening on 30/3/18, some extra days.

Manfred asks about re-cabling of BPMs - This will be discussed. 7 sectors baked.